OpenStudy (bellezabrilla):

Joan's bacteria project produced 8.2x10^8 cells and Kim's project produced 5.8x10^8 cells. The girls determined that Joan's project produced 2.4x10^8 more cells than Kim's project. What is 2.4 x10^8 in expanded notation ? PLZZ HELP !!!

OpenStudy (anonymous):

2.4x10^0=2.4 2.4x10^1=24.0 2.4x10^2=240.0 2.4x10^3=2400.0 Carry on until you get to the eighth power. The number in the power moves the decimal place by that number of places.

OpenStudy (bellezabrilla):

@tom982 I did that and I got 240,000,000

OpenStudy (anonymous):

Yep, that's the answer :)